OpenStudy (anonymous):

The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 70 cm when the weight is 3 kg, what is the distance when the weight is 7 kg? The distance is ? cm

OpenStudy (anonymous):

If the distance is \(x\), the n \(3x=7\times 70 \implies 3x=490 \implies x=\frac{490}{3}\text{ cm}.\)

OpenStudy (anonymous):

A different way to look at it: d=kw (distance (k) varies (=) directly (k) as weight (w) 70=k(3), so k=(70/3) d=(70/3)(7)=(490/3)

OpenStudy (anonymous):

Sorry, I lost my internet connection.

OpenStudy (anonymous):

I have a really hard time when it comes to cm, kg, etc. Because I am an American, I can only see things in feet, inches, etc. This really confuses me.

OpenStudy (anonymous):

Okay Anwar...I DID get 490/3...but when you divided 490 by 3 you get 163.33333, etc, which would be 164 cm. This is the answer I put and it said it was not correct. Now 163 3/3 would not be correct either because 3/3 = 1. Where am I going wrong here?

OpenStudy (anonymous):

Oh wait, no it wouldn 't be 164. It would be 162 and 1/3....maybe that is where I am making my mistake?

OpenStudy (anonymous):

I mean 163 and 1/3

OpenStudy (anonymous):

ahhh...I think I am just confused.

OpenStudy (anonymous):

163 and 1/3 would be correct, unless it is asking for the nearest cm, then it is just 163.

OpenStudy (anonymous):

hmm, let me try that.

OpenStudy (anonymous):

HAHAHA... 163 1/3 was correct. Thank you!!